OpenStudy (anonymous):

7(s-5)=42 how do I do this someone please help and explain it to me

OpenStudy (anonymous):

Are you trying to solve for s?

OpenStudy (anonymous):

yes I am

OpenStudy (anonymous):

s=11

OpenStudy (anonymous):

7s-35=42

OpenStudy (anonymous):

yes I know that s =11 cause they give me the answer but I have to explain and idk how to

OpenStudy (anonymous):

7s=35+42 (7s=77) (s=77/7) (s=11)

OpenStudy (anonymous):

First you divide by 7 on both sides of the equal sign [7(s-5)]/7=42/7 Next you move the 5 to the other side of the equal sign S=6+5 S=11

OpenStudy (anonymous):

\[7(s-5)=42 \] now divide both sides by 7 \[s-5=6\] now add 5 to both sides \[s=11\]

OpenStudy (anonymous):

PEMDAS also will help.

OpenStudy (anonymous):

oh thank all of you guys

OpenStudy (anonymous):

may someone help me please !!! Evaluate 5|x3 - 2| + 7 when x = -2